THERMO Question 1 A pump steadily delivers 13.7 kg/s of water at the conditions given below....

THERMO

Question 1

  1. A pump steadily delivers 13.7 kg/s of water at the conditions given below. Calculate the pump power (hp).

    There is no heat transfer from the pump to the surroundings (i.e., adiabatic conditions).

    Pump Inlet Temperature = 20oC

    Pump Inlet Pressure = 170 kPa

    Pump Inlet Diameter = 11.4 cm

    Pump Inlet Elevation = 20 m

    Pump Exit Temperature = 20oC

    Pump Exit Pressure = 581 kPa

    Pump Exit Diameter = 5 cm

    Pump Exit Elevation = 20 m

Question 2

  1. A well-insulated, full, water tank containing 123 kg of water initially at 44.7oC is being supplied with 24.3oC water at a rate of 4 kg/min. Water is exiting the tank at the same flow rate so that the mass of water in the tank remains constant. Circulation in the tank keeps the water at a uniform temperature. The specific heat of the water is given as C = 4.2 kJ/kg-K. Determine the temperature of the tank (oC) after 6.6 minutes.

    Note: The energy of water in the tank is given as: E = mcv*C*T

    The solution for an equation of the form: a*dT/dt + T + b = 0     is   

    T(t) = k1*exp(-t/a) + k2

    For this problem, the given constraints are:

    T(t=0) = Tinitial

    T(t=infinity) = Tinlet

    Hint: Use the given constraints to determine the constants k1 and k2
Answer #1

See the attached images for the solution.

Question 1 Specific enthal and dens汁 o( a20°c and noKPa hl= 34-078 kJ/Kg 8425am ve ivenloi da rS masS water e Knaus hal13. 6. 388m/s 4 flow enemy 2 2 once wosk is done n the system, so hal Used oth Wi-2 Now we ves 1-2 w,-,- wook clone the pump2, 2 ( g4. 464-84.049) X103 +「6-388)-じ344 2 m W,2-13-7 X 40.3.53= 5610, 33W Once 745-7 watts = 1 h ve as 745-7